# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

### v3

@@ C.2 Distance Metric Based on Data Prediction
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N-1}$, trained diffusion model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$, number of timesteps for simulating the forward process $T$, number of samples used for calculating the mean $S$, maximum and minimum values of $\alpha$ $(\alpha_{s},\alpha_{e})$
0:  $LinearInterp1d(\{\alpha_{t}\}_{t=0,...,T},\{v_{t}\}_{t=0,...,T})$
1:  # Compute $\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right]$
2:  $\overline{D^{2}_{t}}\leftarrow 0,\ \forall{t}\in\{0,1,...,T\}$
3:  $\Delta\alpha\leftarrow\frac{1}{T}(\alpha_{s}-\alpha_{e})$
4:  for $n$ in $randperm(N)[:S]$ do
5:     $\bm{x}_{0}\leftarrow\bm{x}^{(n)}$
6:     if latent-space diffusion model then
7:        $\bm{x}_{0}\leftarrow autoencoder(\bm{x}_{0})$
8:     end if
9:     $\bm{y}_{0}\leftarrow\bm{x}_{0}$
10:
11:     # Simulate the forward process
12:     for $t=1$ to $T$ do
13:        $\alpha_{t}\leftarrow\alpha_{s}-\Delta\alpha t,\ \sigma_{t}\leftarrow\sqrt{1-\alpha_{t}^{2}}$
14:        $\beta_{t}\leftarrow\frac{\alpha_{t}}{\alpha_{t-1}},\ \delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
15:        $\bm{x}_{t}\leftarrow\beta_{t}\bm{x}_{t-1}+\delta_{t}\times\mathcal{N}(\bm{0},\bm{I})$
@@
16:        $\bm{y}_{t}\leftarrow\frac{1}{\alpha_{t}}\left(\bm{x}_{t}-\sigma_{t}\bm{\varepsilon}_{\theta}(\bm{x}_{t},\alpha_{t})\right)$
17:        $\overline{D^{2}_{t}}\leftarrow\overline{D^{2}_{t}}+\frac{1}{S}\left\|\bm{y}_{t}-\bm{y}_{t-1}\right\|_{2}^{2}$
18:     end for
19:  end for
20:
21:  # Compute $v(\alpha)$ using $\overline{D^{2}_{t}}$
22:  for $t=1$ to $T$ do
23:     $v_{t}\leftarrow\sqrt{\frac{1}{\Delta\alpha}\overline{D^{2}_{t}}}$
24:  end for
25:  $v_{0}\leftarrow v_{1}$
CRS-$v_{x}$ can be used for optimizing the training schedule as well. A pseudocode for training diffusion models with CRS is given in Algorithm 3. Since the data prediction is expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$, $v_{x}(\alpha)$ tends to take extremely large values when $\alpha\simeq 0$. To avoid this, $v_{x}(\alpha)$ is evaluated within the range of $\alpha_{\mathrm{th}}\leq\alpha\leq\alpha_{\max}$, and we set $v_{x}(\alpha<\alpha_{\mathrm{th}})=v_{x}(\alpha_{\mathrm{th}})$. Similar to the implementation of the adaptive noise schedule in VDM++ [13], the range of $\alpha$ from $\alpha_{\mathrm{th}}$ to $\alpha_{\max}$ is divided into $B$ evenly spaced bins, and $\overline{D^{2}_{x}}(\alpha,\alpha+\Delta\alpha)$ is approximately evaluated using the exponential moving average (EMA). We used a common configuration of hyperparameters ($B=100,e=0.995,\alpha_{\max}=1,\alpha_{\min}=0,\alpha_{\mathrm{th}}=0.01,\Delta\alpha=10^{-3},\xi=1.0)$ in this study. The initial noise schedule is set to $\alpha(t)=\alpha_{\max}-(\alpha_{\max}-\alpha_{\min})\frac{t}{T}$. Although not explicitly included in the pseudocode, the update of $\overline{D^{2}_{b}}$ starts after 1000 iterations, and the noise schedule $\alpha(t)$ is updated every 100 iterations.
The training times of the diffusion models are listed in Table 5. When using CRS for optimizing training schedule, the training time increases by 20–30% compared with conventional noise schedules because CRS requires evaluating $v_{x}(\alpha)$ during training. To mitigate the increase in training time, one potential approach is to fix the training schedule after a certain number of epochs. In practice, CRS-$v_{x}$ showed little change beyond 10 epochs on LSUN Horse 256 $\times$ 256, suggesting the possibility of reducing training time while maintaining performance by fixing the training schedule after 10 epochs. Reducing the update frequency of $\overline{D^{2}_{x}}$ can be also an effective strategy. Developing an efficient implementation to reduce the training time remains as future work.
Algorithm 3 Adaptive noise schedule for training.
0:  training dataset $\{\bm{x}^{(n)}\}_{n=0,...,N}$, initial noise schedule $\alpha(t)$, number of bins $B$, decay rate of EMA $e$, maximum and minimum values of $\alpha$ for training schedule $(\alpha_{\max},\alpha_{\min})$, infinitesimal variation of $\alpha$ for evaluating the derivative $\Delta\alpha$, minimum $\alpha$ for evaluating $v(\alpha)$ $\alpha_{\mathrm{th}}$, hyperparameter of CRS $\xi$
0:  noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$
1:  $\alpha_{b}\leftarrow\alpha_{\mathrm{th}}+(\alpha_{\max}-\alpha_{\mathrm{th}})\frac{b}{B},\ \forall{b}\in\{0,...,B\}$
2:  $\overline{D^{2}_{b}}\leftarrow 1e^{-6},\ \forall{b}\in\{0,...,B-1\}$
3:  for $n$ in $randperm(N)$ do
4:     # Compute loss for training diffusion models
5:     $\bm{x}_{0}\leftarrow\bm{x}^{(n)},\ t\sim\mathcal{U}(0,1),\ \bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
6:     $\alpha\leftarrow\alpha(t),\ \sigma\leftarrow\sqrt{1-\alpha^{2}}$
7:     $\bm{\varepsilon}_{\alpha}\leftarrow\bm{\varepsilon}_{\theta}(\alpha\bm{x}_{0}+\sigma\bm{\varepsilon},\alpha)$
8:     $loss\leftarrow\frac{1}{2}\left\|\bm{\varepsilon}_{\alpha}-\bm{\varepsilon}\right\|_{2}^{2}$
9:
10:     # Evaluate $\overline{D^{2}_{x}}$ using EMA (without gradient)
11:     if $\alpha\geq\alpha_{\mathrm{th}}$ then
12:        $\alpha^{\prime}\leftarrow\alpha-\Delta\alpha,\ \beta^{\prime}\leftarrow\frac{\alpha^{\prime}}{\alpha},\ \delta^{\prime}\leftarrow\sqrt{1-\beta^{\prime 2}}$
13:        $\bm{x}_{\alpha^{\prime}}\leftarrow\beta^{\prime}\bm{x}_{\alpha}+\delta^{\prime}\times\mathcal{N}(\bm{0},\bm{I})$
14:        $\bm{\varepsilon}_{\alpha^{\prime}}\leftarrow\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})$
15:        $\bm{y}_{\alpha}\leftarrow\frac{1}{\alpha}\left(\bm{x}_{\alpha}-\sigma\bm{\varepsilon}_{\alpha}\right),\ \bm{y}_{\alpha^{\prime}}\leftarrow\frac{1}{\alpha^{\prime}}\left(\bm{x}_{\alpha^{\prime}}-\sigma^{\prime}\bm{\varepsilon}_{\alpha^{\prime}}\right)$
16:        $b\leftarrow\min\left(\left\lfloor\frac{\alpha-\alpha_{\mathrm{th}}}{\alpha_{\max}-\alpha_{\mathrm{th}}}B\right\rfloor,B-1\right)$
17:        $\overline{D^{2}_{b}}\leftarrow e\overline{D^{2}_{b}}+(1-e)\left\|\bm{y}_{\alpha}-\bm{y}_{\alpha^{\prime}}\right\|_{2}^{2}$
18:        $v_{b}\leftarrow\sqrt{\frac{1}{\Delta\alpha}\overline{D^{2}_{b}}}$
19:     end if
20:
21:     # Update training schedule $\alpha(t)$
22:     $\mathcal{X}\leftarrow\{\alpha_{\min},\alpha_{0},\alpha_{1},...,\alpha_{B}\}$
23:     $\mathcal{Y}\leftarrow\{v_{0},v_{0},v_{1},...,v_{B-1},v_{B-1}\}$
24:     $v(\alpha)\leftarrow LinearInterp1d(\mathcal{X},\mathcal{Y})$
25:     Update $\alpha(t)$ and model parameters $\theta$
@@
Table 5: Training time of diffusion models.
Diffusion type Dataset GPUs Epochs Training schedule Training time
Latent space LSUN Church A100 $\times$ 8 1000
Linear
VDM++
CRS-$v_{\mathrm{FID}}$
85 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
105 hours
LSUN Bedroom A100 $\times$ 8 100
CRS-$v_{\mathrm{FID}}$
190 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
245 hours
ImageNet A100 $\times$ 8 500
Linear
VDM++
CRS-$v_{\mathrm{FID}}$
430 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
525 hours
Pixel space LSUN Horse H100 $\times$ 8 50
Linear
Shifted cosine
VDM++
170 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
215 hours
LSUN Bedroom H100 $\times$ 8 50
CRS-$v_{x}$
325 hours
Appendix D Hyperparameter Tuning for Combining Multiple Distance Metrics
CRS-$v_{x}+v_{\mathrm{FID}}$ has four hyperparameters; $w_{x}$, $w_{\mathrm{FID}}$, $\xi_{x}$, and $\xi_{\mathrm{FID}}$. We tuned these hyperparameters in two steps. In the first step, we fixed $\xi_{x}=\xi_{\mathrm{FID}}=1$ and searched for appropriate values of $w_{x}$ and $w_{\mathrm{FID}}$, subject to the constraint $w_{x}+w_{\mathrm{FID}}=1$. In the second step, we fixed $w_{x}$ and $w_{\mathrm{FID}}$, then adjusted $\xi_{x}$ and $\xi_{\mathrm{FID}}$.
@@ Appendix D Hyperparameter Tuning for Combining Multiple Distance Metrics
The detailed results on LSUN Horse 256 $\times$ 256 are listed in Table 6. In the first step, adjusting $w_{x}$ and $w_{\mathrm{FID}}$ led to significant changes in FID at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$. In the second step, the improvement from tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ was small at $\mathrm{NFE}=250$. In contrast, for $\mathrm{NFE}=50$, the tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ resulted in a substantial improvement in FID, indicating that the importance of hyperparameter tuning increases as NFE decreases.
@@
Table 6: Detailed results of hyperparameter tuning on LSUN Horse 256 $\times$ 256 in pixel-space diffusion model. DPM-Solver++(2M) was used as sampler. To conserve computational resources, FID score was calculated using 10K samples.
@@
The dependence of noise schedules on the hyperparameters is shown in Figs. 7, 7, and 7. Increasing either $w_{\mathrm{FID}}$ or $\xi_{\mathrm{FID}}$ amplifies the impact of $v_{\mathrm{FID}}(\alpha)$, resulting in more timesteps being allocated to the region where $\alpha\simeq 1$ (see Figs. 7 and 7). On the other hand, increasing either $w_{x}$ or $\xi_{x}$ strengthens the impact of $v_{x}(\alpha)$, causing more timesteps to be allocated to the region where $\alpha\simeq 0$ (see Figs. 7 and 7).
The hyperparameter settings for CRS-$v_{x}+v_{\mathrm{FID}}$ are listed in Table 7. Noise schedules $\alpha(t)$ obtained using CRS-$v_{x}+v_{\mathrm{FID}}$ are illustrated in Figs. 13, 13, 13, 13, 13, and 13.
Table 7: Hyperparameter settings of CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse 256 $\times$ 256, LSUN Bedroom 256 $\times$ 256, FFHQ 64 $\times$ 64, and CIFAR10 32 $\times$ 32.
@@
LSUN Horse 256 $\times$ 256 SDE-DPM-Solver++(2M) 250 0.9 0.1 1.4 1.0
@@
LSUN Bedroom 256 $\times$ 256 DPM-Solver++(2M) 10 0.5 0.5 1.0 1.0
@@
FFHQ 64 $\times$ 64 DPM-Solver++(2M) 50 0.5 0.5 1.6 1.4
@@
10 0.5 0.5 1.6 1.5
iPNDM 50 0.5 0.5 1.2 1.5
20 0.5 0.5 1.4 1.4
10 0.5 0.5 1.5 1.4
CIFAR10 32 $\times$ 32 DPM-Solver++(2M) 50 0.5 0.5 1.6 1.0
20 0.5 0.5 1.8 2.4
10 0.5 0.5 1.6 3.0
iPNDM 50 0.7 0.3 1.6 1.2
@@
10 0.5 0.5 1.6 3.2
Figure 8: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with SDE-DPM-Solver++(2M).
Figure 9: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with PNDM.
Figure 10: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Horse 256 $\times$ 256 with DPM-Solver++(2M).
Figure 11: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Bedroom 256 $\times$ 256 with DPM-Solver++(2M).
Figure 12: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating LSUN Bedroom 256 $\times$ 256 with UniPC.
Figure 13: CRS-$v_{x}+v_{\mathrm{FID}}$ based noise schedules used for generating FFHQ 64 $\times$ 64 with DPM-Solver++(2M).

### v4

@@ C.2 Discrepancy Measure Based on Data Prediction
1:training dataset $\{\bm{x}^{(n)}\}_{n=0,\dots,N-1}$, trained diffusion model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$, number of timesteps $T$, number of samples for mean $S$, range of $\alpha$ $(\alpha_{s},\alpha_{e})$
[…]
3:# Evaluate $\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\!\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right]$
[…]
6:for $n$ in $\mathrm{randperm}(N)[:S]$ do
7:  $\bm{x}_{0}\leftarrow\bm{x}^{(n)}$
8:  if model operates in latent space then
9:   $\bm{x}_{0}\leftarrow\mathrm{autoencoder}(\bm{x}_{0})$
10:  end if
11:  $\bm{y}_{0}\leftarrow\bm{x}_{0}$
12:
13:  # Simulate the forward process
14:  for t=1t=1toTT do
15:   $\alpha_{t}\leftarrow\alpha_{s}-\Delta\alpha\,t,\quad\sigma_{t}\leftarrow\sqrt{1-\alpha_{t}^{2}}$
16:   $\beta_{t}\leftarrow\dfrac{\alpha_{t}}{\alpha_{t-1}},\quad\delta_{t}\leftarrow\sqrt{1-\beta_{t}^{2}}$
17:   $\bm{x}_{t}\leftarrow\beta_{t}\bm{x}_{t-1}+\delta_{t}\,\mathcal{N}(\bm{0},\bm{I})$
@@
18:   $\bm{y}_{t}\leftarrow\dfrac{1}{\alpha_{t}}\!\left(\bm{x}_{t}-\sigma_{t}\,\bm{\varepsilon}_{\theta}(\bm{x}_{t},\alpha_{t})\right)$
19:   $\overline{D^{2}_{t}}\leftarrow\overline{D^{2}_{t}}+\dfrac{1}{S}\left\|\bm{y}_{t}-\bm{y}_{t-1}\right\|_{2}^{2}$
20:  end for
21:end for
22:
23:# Compute $v(\alpha)$ using $\overline{D^{2}_{t}}$
24:for t=1t=1toTT do
25:  $v_{t}\leftarrow\sqrt{\dfrac{1}{\Delta\alpha}\,\overline{D^{2}_{t}}}$
@@
[…]
Table 9: Hyperparameters used to compute $v_{x}(\alpha)$ for optimizing sampling schedule.
Parameter Description Used Value
$T$ Number of timesteps for simulating the forward process $10^{3}$
$S$ Number of samples used to estimate $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ $10^{4}$
$\alpha_{s}$ Maximum $\alpha$ used to evaluate $v(\alpha)$ 1.0
$\alpha_{e}$ Minimum $\alpha$ used to evaluate $v(\alpha)$ 0.0
Algorithm 3 shows how $v_{x}(\alpha)$ can be used to adaptively optimize the training schedule. Since the data prediction is expressed as $\bm{x}_{\theta}(\bm{x},\alpha)=\frac{\bm{x}-\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)}{\alpha}$, $v_{x}(\alpha)$ can become numerically unstable as $\alpha\to 0$. To prevent this, we restrict evaluation to the range $\alpha_{\mathrm{th}}\leq\alpha\leq\alpha_{\max}$ and fix $v_{x}(\alpha)=v_{x}(\alpha_{\mathrm{th}})$ for $\alpha<\alpha_{\mathrm{th}}$. Similar to VDM++ (Kingma and Gao, 2023), we divide the range of $\alpha$ into $B$ uniform bins and estimate $\overline{D^{2}_{x}}$ for each bin using an exponential moving average (EMA).
Hyperparameters are summarized in Table 10. We use the same configuration for all our experiments, regardless of the diffusion type or dataset. Although not explicitly included in the pseudocode, the update of $\overline{D^{2}_{b}}$ starts after 1000 iterations, and the noise schedule $\alpha(t)$ is updated every 100 iterations.
While this adaptive scheme slightly increases training time (by 20–30% compared to fixed schedules; see Table 15 in Appendix E), we find that the schedule stabilizes early: for example, on LSUN Horse $\times256$, $v_{x}(\alpha)$ changes very little after 10 epochs. Thus, fixing the schedule after a few epochs or reducing the frequency of $\overline{D^{2}_{x}}$ updates may mitigate the cost without degrading performance. Developing a more efficient implementation remains an important direction for future work.
Algorithm 3 Adaptive optimization of training schedule using $v_{x}(\alpha)$.
1:training dataset $\{\bm{x}^{(n)}\}_{n=0,\dots,N}$, initial noise schedule $\alpha(t)$, number of bins $B$, decay rate of EMA $e$, range of $\alpha$ $(\alpha_{\max},\alpha_{\min})$, derivative step $\Delta\alpha$, minimum value for evaluating $v(\alpha)$ $\alpha_{\mathrm{th}}$, hyperparameter of CRS $\xi$
2:noise prediction model $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$
[…]
5:for $n$ in $\mathrm{randperm}(N)$ do
6:  # Compute loss for training diffusion models
7:  $\bm{x}_{0}\leftarrow\bm{x}^{(n)},\quad t\sim\mathcal{U}(0,1),\quad\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
8:  $\alpha\leftarrow\alpha(t),\quad\sigma\leftarrow\sqrt{1-\alpha^{2}}$
9:  $\bm{\varepsilon}_{\alpha}\leftarrow\bm{\varepsilon}_{\theta}(\alpha\bm{x}_{0}+\sigma\bm{\varepsilon},\,\alpha)$
10:  $loss\leftarrow\dfrac{1}{2}\left\|\bm{\varepsilon}_{\alpha}-\bm{\varepsilon}\right\|_{2}^{2}$
11:
12:  # Evaluate $\overline{D^{2}_{b}}$ using EMA (without gradient)
13:  if $\alpha\geq\alpha_{\mathrm{th}}$ then
14:   $\alpha^{\prime}\leftarrow\alpha-\Delta\alpha,\quad\sigma^{\prime}\leftarrow\sqrt{1-\alpha^{\prime 2}},\quad\beta^{\prime}\leftarrow\dfrac{\alpha^{\prime}}{\alpha},\quad\delta^{\prime}\leftarrow\sqrt{1-{\beta^{\prime}}^{2}}$
15:   $\bm{x}_{\alpha^{\prime}}\leftarrow\beta^{\prime}\bm{x}_{\alpha}+\delta^{\prime}\mathcal{N}(\bm{0},\bm{I})$
16:   $\bm{\varepsilon}_{\alpha^{\prime}}\leftarrow\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\,\alpha^{\prime})$
17:   $\bm{y}_{\alpha}\leftarrow\dfrac{1}{\alpha}\!\left(\bm{x}_{\alpha}-\sigma\bm{\varepsilon}_{\alpha}\right),\quad\bm{y}_{\alpha^{\prime}}\leftarrow\dfrac{1}{\alpha^{\prime}}\!\left(\bm{x}_{\alpha^{\prime}}-\sigma^{\prime}\bm{\varepsilon}_{\alpha^{\prime}}\right)$
18:   $b\leftarrow\min\!\left(\left\lfloor\dfrac{\alpha-\alpha_{\mathrm{th}}}{\alpha_{\max}-\alpha_{\mathrm{th}}}\,B\right\rfloor,\,B-1\right)$
19:   $\overline{D^{2}_{b}}\leftarrow e\,\overline{D^{2}_{b}}+(1-e)\left\|\bm{y}_{\alpha}-\bm{y}_{\alpha^{\prime}}\right\|_{2}^{2}$
20:   $v_{b}\leftarrow\sqrt{\dfrac{1}{\Delta\alpha}\,\overline{D^{2}_{b}}}$
21:  end if
22:
23:  # Update training schedule $\alpha(t)$
24:  $\mathcal{X}\leftarrow\{\alpha_{\min},\alpha_{0},\alpha_{1},\dots,\alpha_{B}\}$
25:  $\mathcal{Y}\leftarrow\{v_{0},v_{0},v_{1},\dots,v_{B-1},v_{B-1}\}$
26:  $v(\alpha)\leftarrow\mathrm{LinearInterp1d}(\mathcal{X},\mathcal{Y})$
27:  Update noise schedule $\alpha(t)$ by solving Eq. (9)
28:
29:  # Update model parameters $\theta$
30:  loss.backward()
31:  optimizer.step()
32:end for
Table 10: Hyperparameters used for optimizing the training schedule with CRS-$v_{x}$.
Parameter Description Used Value
$\alpha_{\min}$ Minimum value of $\alpha$ for the training schedule 0.0
$\alpha_{\max}$ Maximum value of $\alpha$ for the training schedule 1.0
$\Delta\alpha$ Small step used for approximating the derivative $10^{-3}$
$\alpha_{\mathrm{th}}$ Threshold of $\alpha$ for clipping $v(\alpha)$ 0.01
$\xi$ A parameter controlling the dependence of $\alpha(t)$ on $v(\alpha)$ 1
$B$ Number of bins used to divide the range of $\alpha$ 100
$e$ EMA decay rate for approximately evaluating $\overline{D_{x}^{2}}(\alpha,\alpha^{\prime})$ 0.995
$\alpha(t)$ Initial noise schedule 1t1-t
Appendix D Hyperparameter Tuning for Combining Multiple Discrepancy Measures
CRS-$v_{x}+v_{\mathrm{FID}}$ achieves the best performance for sampling in pixel-space diffusion models. CRS-$v_{x}+v_{\mathrm{FID}}$ involves four hyperparameters: $w_{x}$, $w_{\mathrm{FID}}$, $\xi_{x}$, and $\xi_{\mathrm{FID}}$. The effects of these hyperparameters on the resulting noise schedules are illustrated in Figs. 6, 6, and 6. Increasing either $w_{\mathrm{FID}}$ or $\xi_{\mathrm{FID}}$ amplifies the influence of $v_{\mathrm{FID}}(\alpha)$, resulting in more timesteps being allocated to the region where $\alpha\simeq 1$ (see Figs. 6 and 6). Conversely, increasing either $w_{x}$ or $\xi_{x}$ strengthens the contribution of $v_{x}(\alpha)$, leading to more timesteps being assigned to the region where $\alpha\simeq 0$ (see Figs. 6 and 6).
Figure 4: Effect of the weights $w_{x}$ and $w_{\mathrm{FID}}$ on the noise schedule generated by CRS-$v_{x}+v_{\mathrm{FID}}$, with fixed exponents $\xi_{x}=\xi_{\mathrm{FID}}=1$.
Figure 5: Effect of the exponent $\xi_{\mathrm{FID}}$ on the noise schedule generated by CRS-$v_{x}+v_{\mathrm{FID}}$, with fixed weights $w_{x}=w_{\mathrm{FID}}=0.5$.
Figure 6: Effect of the exponent $\xi_{x}$ on the noise schedule generated by CRS-$v_{x}+v_{\mathrm{FID}}$, with fixed weights $w_{x}=w_{\mathrm{FID}}=0.5$.
We tuned these hyperparameters in two steps. First, we fixed $\xi_{x}=\xi_{\mathrm{FID}}=1$ and searched for appropriate values of $w_{x}$ and $w_{\mathrm{FID}}$ under the constraint $w_{x}+w_{\mathrm{FID}}=1$. Then, we fixed the weights and adjusted $\xi_{x}$ and $\xi_{\mathrm{FID}}$. Although further iterations between tuning weights and exponents are possible, we found that just one iteration per step was typically sufficient to outperform conventional noise schedules.
@@ Appendix D Hyperparameter Tuning for Combining Multiple Discrepancy Measures
The detailed tuning results on LSUN Horse $\times256$ are presented in Table 11. In the first step, adjusting $w_{x}$ and $w_{\mathrm{FID}}$ led to substantial changes in FID at both $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$. In the second step, tuning $\xi_{x}$ and $\xi_{\mathrm{FID}}$ had little impact at $\mathrm{NFE}=250$, but resulted in a significant FID improvement at $\mathrm{NFE}=50$. This suggests that hyperparameter sensitivity increases as the NFEs decrease.
@@
Table 11: Detailed results of hyperparameter tuning for CRS-$v_{x}+v_{\mathrm{FID}}$ on LSUN Horse $\times256$ in pixel-space diffusion model. DPM-Solver++(2M) was used as sampler. To reduce computational cost, FID scores were calculated using 10K samples.
@@
The final hyperparameter settings used in our experiments are summarized in Table 12. The resulting noise schedules $\alpha(t)$ for CRS-$v_{x}+v_{\mathrm{FID}}$ are visualized in Figs. 11, 11, 11, 11, and 11.
Table 12: Hyperparameter settings used for CRS-$v_{x}+v_{\mathrm{FID}}$ in our experiments on LSUN Horse $\times256$, LSUN Bedroom $\times256$, FFHQ $\times64$, and CIFAR10 $\times32$.
@@
LSUN Horse $\times256$ SDE-DPM-Solver++(2M) 250 0.9 0.1 1.4 1.0
@@
LSUN Bedroom $\times256$ DPM-Solver++(2M) 10 0.5 0.5 1.0 1.0
@@
FFHQ $\times64$ DPM-Solver++(2M) 20 0.5 0.5 1.7 1.4
@@
UniPC(3M) 20 0.5 0.5 1.7 1.4
@@
iPNDM(3M) 20 0.5 0.5 1.6 1.2
10 0.5 0.5 1.6 1.2
iPNDM(4M) 20 0.5 0.5 1.4 1.4
10 0.5 0.5 1.6 1.4
CIFAR10 $\times32$ DPM-Solver++(2M) 20 0.5 0.5 1.8 1.2
10 0.5 0.5 1.7 1.2
UniPC(3M) 20 0.5 0.5 2.3 1.0
10 0.5 0.5 1.9 1.5
iPNDM(3M) 20 0.5 0.5 1.7 1.4
10 0.5 0.5 1.7 1.4
iPNDM(4M) 20 0.5 0.5 1.6 1.6
10 0.5 0.5 1.6 1.6
Figure 7: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse $\times256$, used with SDE-DPM-Solver++(2M) during sampling.
Figure 8: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse $\times256$, used with PNDM during sampling.
Figure 9: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Horse $\times256$, used with DPM-Solver++(2M) during sampling.
Figure 10: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Bedroom $\times256$, used with DPM-Solver++(2M) during sampling.
Figure 11: Noise schedules generated by CRS-$v_{x}+v_{\mathrm{FID}}$ for LSUN Bedroom $\times256$, used with UniPC during sampling.